the length of a rectangle is 10 units longer than its width. if the total perimeter of the rectangle is 44 units, what is the width?
Answer by oberobic(2304)

P = perimeter
L = length
W = width
P = 2L + 2W = 2(L+W)
P = 44
.
L = W+10
.
Substitute
.
2(W+10) + 2W = 44
2W + 20 + 2W = 44
4W = 24
W = 6
.
L = W+10
L = 6+10
L = 16
.
Check your work!
2L = 32
2W = 12
32+12 = P = 44
Correct.
.
Re-read the question to make sure you answer it.
The width is 6.
.
Done.